-4(1-3y)-6Y=-5
------
2

Answer by stanbon(75887)

-4(1-3y)-6Y=-5/2
---------------
-8(1-3y)-12Y = -5
------------------------
-8 + 24y - 12y = -5
-------
12y = 3
y = 1/4
----------------------
Cheers,
Stan H.
------------